Project

General

Profile

NDOS CommissionMtng 20101206

Attendance: Jon, Sue, Jarek, Peter, Mark Messier, Minerba, Nate, Mat, Andrew, Peter, Alena, Kurt, Leon, Susan, Denis, Ryan, Xinjie, Ron, Rick, Greg

Agenda:
  1. DAQ/Electronics Update & Plans
  2. DCS Status & Plans
  3. Shifters' report
  4. AOB

DAQ:

  • Andrew:
    • Updates to the DCM kernel and DcmProgUtils broke the online software "test" release due to changes in "Set" and "Get" methods.
    • Readbacks from FEB "work", but he is seeing non-sensical values
  • Peter: much effort is going on in understanding the trigger timing, which will likely consume the effort over the next couple of days:
    • There appears to be an anomalous trigger offset
    • We observe a 55 us jitter
    • It would be good to add the Booster trigger into the readout (Holger responded that he would add this)
    • It would be good to put a GPS+TDU on top of the MINOS building; unfortunately we don't have a spare TDU at the moment.
  • Sue:
    • Mat & Denis were reporting difficulties getting runs started; mostly failing when (some) DCMs were configuring hardware.
    • To improve the likelihood of a run actually starting, Sue changed the max. number of allowed attempts to configure the hardware on a DCM. That doesn't fix the underlying problem, but should help in the near term.
    • More info is now being printed by the DCMApplication which should be useful for debgugging
    • It appears that error bits are not clearing.
  • Jon: Kurt and Jon are working on getting the DCM hardware configuration into the database

DCS:

  • Chad reported that he wired up the chiller on Friday and the data are being readout in Labview.
  • Plan is to work with Athans this week on getting the DCM DCS data readout

Shifter report:

  • Nothing specific, but documentation is really needed...